### v1

# Globally defined Carroll symmetry of gravitational waves

Abstract
The locally defined Carroll symmetry of a gravitational wave is extended to a globally defined one. Translations and Carroll boosts associated with two independent globally defined solutions of a Sturm-Liouville equation allow us to describe the motions. The Displacement Memory Effect arises for particular choices of the parameters which yield trajectories with zero momentum. We illustrate our general statements by the Pöschl-Teller profile.
pacs:
04.20.-q Classical general relativity
preprint: Global Carroll v1. arXiv: [gr-qc, hep-th]
I Introduction
Braginsky and Thorne had suggested that gravitational waves might be observed by detecting the displacement of particles initially in rest, called later the Memory Effect BraTho . While experimental verification is still in the making, an insight can be gained by taking advantage of the Carroll symmetry Leblond ; SenGupta . Baldwin-Jeffery-Rosen (BJR) coordinates BaJe ; Rosen provide a simple description Sou73 ; Carroll4GW , however suffer of being regular only in finite intervals, requiring to glue them together. In this paper we show that such complications can be avoided by switching to Brinkmann (B) coordinates Bri by solving a matrix Sturm-Liouville equation, (II.3) below. The price to pay is, though, to get more complicated-looking expressions.
We illustrate our general theory by the Pöschl-Teller profile PTeller ; Plyushchay09 ; ChaKar , which is a very good analytical approximation of the widely studied Gaussian one. We follow mostly our review EZHRev which provides the reader also with further references.
Our notations are: $({\bm{X}},U,V)$ are Brinkmann coordinates and $({\bm{x}},u,v)$ are BJR coordinates on relativistic spacetime. $U$ is an affine parameter for geodesics. $\Theta^{Brink}={\bm{\xi}}{\partial}_{{\bm{X}}}+\eta{\partial}_{V}$ denote vector fields which generate the infinitesimal isometries on spacetime.
II From Brinkmann to BJR and back
Plane gravitational waves are conveniently described by using globally defined Brinkmann coordinates Bri ; exactsol in terms of which the metric is,
$g_{\mu\nu}dX^{\mu}dX^{\nu}=d{\bm{X}}^{2}+2dUdV+K_{ij}X^{i}X^{j}\,dU^{2}\,$, (II.1)
where ${\bm{X}}=(X^{1},\dots,X^{D})$ parametrizes the transverse plane, endowed with the flat Euclidean metric $d{\bm{X}}^{2}=\delta_{ij}\,dX^{i}dX^{j}$. $U$ and $V$ are light-cone coordinates and $(K_{ij})$ we call the profile is a symmetric $D\times D$ matrix whose entries depend only on $U$. Henceforth we focus our attention at $D=1$ or $D=2$ and assume, for simplicity, that is diagonal. In $D=2$ we take
$K_{ij}X^{i}X^{j}={\mathcal{A}}(U)\Big((X^{+})^{2}-(X^{-})^{2}\Big)\,$, (II.2)
which yields a vacuum solution of the Einstein equations Bri ; exactsol . We consider “sandwich waves” BoPi89 , whose profile ${\mathcal{A}}(U)$ vanishes outside an interval $U_{i}<U<U_{f}$. The $V$-equation follows from the transverse ones EZHRev and will henceforth not be studied.
After the wave has passed, particles initially at rest exhibit the Velocity Memory Effect (VM) : they move with constant velocity. Under certain “quantization” conditions the outgoing velocity can vanish, though, and we get the Displacement Memory Effect (DM) ZelPol ; DM-1 ; DM-2 ; Sila-PLB , which might well play a rôle in future observations BraTho ; DM-2 .
Plane gravitational waves have long been known to have a $2D+1$ parameter symmetry group composed of $D+1$ translations, completed by $D$ (rather mysterious) transformations BoPiRo ; Sou73 . More recently Carroll4GW , this group was identified with the Carroll group Leblond ; SenGupta , as young Lévy-Leblond called it jokingly. Its algebraic structure is readily determined by switching to Baldwin et al (BJR) coordinates BaJe ; Rosen ; Sou73 which are however only local. This note sheds further light at the Carroll – Memory relation using globally defined Brinkmann coordinates.
We start with considering real solutions of the Sturm-Liouville equations
${P}^{\prime\prime}={\mathcal{A}}(U)P,\qquad(P^{T}){P}^{\prime}=(P^{T})^{\prime}P\,$ (II.3)
for a $D{\times}D$ matrix $P(U)=\big(P_{ij}(U)\big)$ EZHRev . The BJR coordinates $({\bm{x}},u,v)$, are then obtained by
${{\bm{X}}}=P(u){\bm{x}},\qquad U=u,\qquad V=v-\frac{1}{4}{\bm{x}}\cdot\big(P^{T}{}P\big)^{\prime}{\bm{x}}\,$, (II.4)
in terms of which the metric is,
$g_{\mu\nu}dx^{\mu}dx^{\nu}=\big(P(u)^{T}P(u)\big)_{ij}(u)\,dx^{i}dx^{j}+2du\,dv\,.$ (II.5)
Carroll symmetry consists, in BJR coordinates, of “vertical” ($v$) and transverse ($\bm{x}$}) translations with parameter $h$ and $\bf c$}, completed by Carroll boosts with parameter $\bf b$} Sou73 ; Carroll4GW ,
${\bm{x}}\to{\bm{x}}+{\bf c}+S(u)\,{\bf b}\,,\qquad u\to u,\qquad v\to v-{\bf b}\cdot{\bm{x}}-{{\scriptstyle{\frac{1}{2}}}}{\bf b}\cdot{}S(u)\,{\bf b}+h\,$, (II.6)
where the symmetric $2\times 2$ matrix
$S(u)=\int^{u}_{u_{0}}\!\left(P^{T}P\right)^{-1}\!(t)dt$ (II.7)
is referred to as the Souriau matrix Sou73 ; Carroll4GW ; EZHRev . Infinitesimal boosts parametrized with $b_{j}$ are,
$\theta^{BJR}=h\partial_{v}+c^{j}\partial_{j}+b_{j}(S^{ji}\partial_{i}-x^{j}\partial_{v})\,.$ (II.8)
Finding the trajectories is then straightforward. Noether’s theorem provides us with $2D+1$ associated conserved quantities Sou73 ; Carroll4GW ,
${\bm{p}}={\bm{p}}_{0}=(P^{T}P)(u)\,{{\bm{x}}}^{\prime}(u),\qquad{\bf k}={\bf k}_{0}={\bm{x}}(u)-S(u)\,{\bm{p}}\,$, (II.9)
interpreted as conserved linear and boost-momentum, supplemented by $m={u}^{\prime}=1$. Then the conserved quantities determine the geodesic,
${\bm{x}}(u)={\bf k}_{0}+S(u)\,{\bm{p}}_{0}\,.$ (II.10)
The subtlety comes from that the BJR coordinates are defined only in intervals $I_{k}=[u_{k-1},u_{k}]$ between the (mandatory Sou73 ; Carroll4GW ) zeros of the determinant,
$\det P(u_{k})=0\,$, (II.11)
which play a fundamentally important rôle in our investigation, as we shall see. The BJR coordinates are thus only local : both the metric (II.5) and the Souriau matrix $S$ in (II.7) are singular at the junction points $u_{k}$, as will be illustrated in FIG.1 for the Pöschl - Teller profile. The singularity problem will be resolved in sect. III by switching to Brinkmann coordinates.
Now we set $P^{T}=P$ and we consider $D=1$ (when $P$ is a scalar function) for simplicity, although our results and conclusions could be extended to non-diagonal profiles and to any dimension. Let us assume that $u<u_{1}$, the first zero of $\det(P)$. Requiring that our particles be in rest before the wave arrives,
${{\bm{x}}}(-\infty)=x_{0}{\quad\text{\small and}\quad}{{\bm{x}}}^{\prime}(-\infty)=0\,$, (II.12)
imply that the momentum vanishes in $I_{1}$,
${\bm{p}}=0\,.$ (II.13)
The Souriau term is thus switched off from (II.10) and the transverse BJR “trajectory” is merely a fixed point.
${\bm{x}}(u)={\bm{x}}_{0}=\mathop{\rm const}\nolimits\,$, (II.14)
confirming, in BJR coordinates, the “no motion for Carroll” maxim Leblond . In contrast, the Brinkmann trajectory is non-trivial : (II.4) yields,
${\bm{X}}(U)=P(u)\,{\bm{x}}_{0}\,$, (II.15)
with $U=u$. The initial conditions (II.12) then require :
$P(-\infty)=1{\quad\text{\small and}\quad}P^{\prime}(-\infty)=0\,.$ (II.16)
The Brinkmann trajectory, given by the $P$-matrix, may look quite complicated, depending on the Sturm-Liouville solution $P$ of (II.4) EZHRev . For appropriate “quantized” values of the wave parameter the wave zone contains an integer number of half-waves; then $P(u)$ tends to a constant matrix also for $u\to\infty$ : DM is obtained DM-1 ; DM-2 ; Sila-PLB .
We emphasise that the investigations above are a priori valid only before the first zero of $\det{P}$ and should then be restarted until the next zero, and so on.
Another remarkable consequence of (II.15) is that when $\det P(u_{0})=0$ for some $u_{0}$, then all transverse trajectories which start from an $x_{0}\in\mathrm{Ker}\,\big(P(u_{0})\big)$ are focused at the origin, and we get a caustic point BoPi89 ; EZHRev . The question will be further discussed in Sec.IV and illustrated in FIG. 2.
III Carroll symmetry in Brinkmann coordinates
Simple as they are, BJR coordinates (II.4) suffer from being valid only in intervals between subsequent zeros of $P$, and the adjacent expressions should be glued together – which may be laborious. A global approach can conveniently be given instead by using Brinkmann coordinates. For simplicity we restrict our attention at $D=1$ transverse dimension and start with a fixed interval $I_{k}$. A straightforward computation then shows that pulling back the BJR expression (II.6) by using (II.4) [or its infinitesimal form (II.8)] yields the Carroll generators written in Brinkmann coordinates,
${\Theta^{Brink}=h\frac{\partial}{\partial V}+c\left(P\frac{\partial}{\partial X}-P^{\prime}X\frac{\partial}{\partial V}\right)+b\left((PS)\frac{\partial}{\partial X}-(PS)^{\prime}X\frac{\partial}{\partial V}\right)\,,}\;\;$ (III.1)
The coefficients of ${\partial}_{X}$ and of ${\partial}_{V}$ are indeed the the symmetry generators $\xi$ and $\eta$, mentioned in the Introduction. In Sec.IV they will be spelt out explicitly for Pöschl - Teller .
The $P$-matrix and its derivative $P^{\prime}$ generate translations, whereas $PS$ and $(PS)^{\prime}$ generate Carroll boosts GenRelGrav . Their only nonzero commutator yields a vertical translation Blau24 ,
$\Big[\underbrace{P\frac{\partial}{\partial X}-P^{\prime}X\frac{\partial}{\partial V}}_{translation}\ ,\ \underbrace{PS\frac{\partial}{\partial X}-(PS)^{\prime}X\frac{\partial}{\partial V}}_{boost}\Big]=-\partial_{V}\,.$ (III.2)
The symmetry is thus the Heisenberg algebra.
$P$ thus plays a double role : by (II.15) it determines the B trajectory, whereas (III.1) says that it also generates translations. The B trajectory is obtained indeed from $x_{0}=\mathop{\rm const}\nolimits$ by a (generally $U$-dependent) translation.
The Brinkmann form (III.1) has an important advantage w.r.t. the BJR expression in (II.6) : multiplication by $P$ supresses the singularity of Souriau matrix $S$ Sou73 ; EZHRev  :
Q(U)=P(U)S(U)Q(U)=P(U)S(U)\, (III.3)
is regular for all $U$. Thus (III.1) extends the Carroll symmetry (II.6) naturally from a fixed interval $I_{k}$ to the entire $U$-axis.
Both $P$ and $Q$ are solutions of the Sturm-Liouville equation (II.3) EZHRev ; Blau24 as it can be verified by a direct calculation. The Wronskian is $W(P,Q)=P^{\prime}Q-Q^{\prime}P=-1$, therefore $P$ and $Q$ are independent. The expression (III.1) could also be verified independently, by checking the symmetry equation $L_{Y}g=0$ Torre ; GenRelGrav .
Contracting the Brinkmann metric (II.1) with (III.1) provides us with,
${\bm{p}}_{0}=P{\bm{X}}^{\prime}-\big(P\big)^{\prime}{\bm{X}}{\quad\text{\small and}\quad}{\bf k}_{0}=-Q{\bm{X}}^{\prime}+\big(Q\big)^{\prime}{\bm{X}}\,$, (III.4)
where ${\bm{p}}_{0}$ and ${\bf k}_{0}$ are, a priori, new constants of the motion. However expressing them in BJR terms, (II.4), the previous expressions (II.9) are recovered; as anticipated by our notation.
The Sturm-Liouville solutions $P$ and $Q$ play also yet another rôle, though : either combining (II.10) with (II.4) or solving (III.4) directly yields the general Brinkmann trajectory,
${{\bm{X}}(U)=P(U)\,{\bf k}_{0}+Q\,{\bm{p}}_{0}\,.}\;\;$ (III.5)
which is our second main result. When ${\bm{p}}_{0}=0$ cf. (II.13), then the $Q$ - term is switched off and we recover (II.15) with ${\bm{x}}_{0}={\bf k}_{0}$. For parameters which correspond to having an integer number of half-waves in the wavezone, we have also $P(+\infty)=\mathop{\rm const}\nolimits$ and we get DM.
IV Illustration by the Pöschl - Teller profile
We illustrate our general theory by the Pöschl - Teller profile PTeller ; Plyushchay09 ; ChaKar ; DM-1 in $D=1$,
${\mathcal{A}}\equiv{\mathcal{A}}^{PT}(U)=-\dfrac{4m(m+1)}{\cosh^{2}U}\,.$ (IV.1)
The Sturm-Liouville equation (II.3) is solved by Legendre functions. DM trajectories arise when $m$ is an integer, and we get Legendre polynomials, $(-1)^{m}P_{m}(\tanh U)$ DM-1 which have $m$ zeros.
Turning to BJR by (II.4), the Souriau matrix $S$ in (II.7) is regular between two subsequent zeros of $P$ but diverges at the junction points. For $m=1$, for example,
$S_{m=1}(u)=u-\coth u\,$ (IV.2)
obtained for the choice $u^{\pm}_{0}\approx\pm 1.2$, is regular either in $I_{-}<0$ or in $I_{-}>0$, but diverges at $u=0$, as depicted in FIG.1a. Similarly for $m=2$,
$S_{m=2}=\frac{1}{4}\left(u+\frac{3\sinh\left(2u\right)}{2(2-\cosh(2u))}\right)\,$, (IV.3)
is singular where the denominator vanishes, as shown in FIG.1b.
Figure 1: The Sturm-Liouville solutions $\bf P$ in (II.3) are DM trajectories written in Brinkmann coordinates with wave numbers $\bf m=1$} and $\bf m=2$}, respectively. The Souriau matrix S diverges where $\bf P=0$, but the pull-back to Brinkmann $\color[rgb]{0,0.5,0}\bf Q=PS$} is regular for all $U$.
A second solution of the Sturm-Liouville equation (II.3), $Q_{m}=P_{m}S_{m}$ , is regular for all $U$ but it is a Legendre function with non-DM behavior. The solutions $P_{m}$ and $Q_{m}$ are independent when the Wronskian does not vanish, $W(P,Q)\neq 0$. For $m=1$ and $2$ we have, for example,
$\begin{array}[]{llclllc}P_{1}(U)&=&{-}\tanh U\,,&\qquad&P_{2}(U)&=&\frac{1}{2}(3\tanh^{2}U-1)\,,\\[6.0pt] Q_{1}(U)&=&{1-U\tanh U},&\qquad&Q_{2}(U)&=&\frac{1}{2}\big(2U-3\tanh U-3U\operatorname{sech}^{2}U\big)\,.\end{array}$ (IV.4)
The two SturmLiouville{Sturm-Liouville\;} solutions are shown in FIG. 1 in red and in green.
The $P(U)$ found above provide us with DM geodesics which sare consistent with (II.15). Their caustic behavior at the zeros of $P$ is manifest in FIG.2.
Figure 2: For a DM wave in $D=1$ transverse dimension Brinkmann trajectories focus to the caustic points determined by $P_{m}(U_{k})=0$, $k=1,\dots,m$. For $m=1$ there is one focal point at the origin, and for $m=2$, we have two of them, determined by $\tanh U_{\pm}=\pm{1}/{\sqrt{3}}$.
Coming to symmetries of Pöschl - Teller , the coefficients of the Carroll generators $P$ and $Q$ in (III.1) are shown in FIGs. 3 and 4.
(a)                       (b)
Figure 3: For wave number $\bf m=1$}, the (a) translation generator ${\bf P}_{1}$ [which is also a trajectory] is odd. (b) the generator of a Carroll boost, ${\bf Q}_{1}={\bf P_{1}S_{1}}$, is even. The only zero of ${\bf P}_{1}$ is at $U=0$.
(a)                       (b)
Figure 4: Infinitesmal Carroll generators (IV) with wave number $\bf m=2$}. (a) For a translation, the transverse component ${\bf P}_{2}$ is even, and has two zeros, at $U_{\mp}$. (b) For a boost, ${\bf Q}_{2}={\bf P_{2}S_{2}}$ is odd.
The globally defined Carroll symmetry generators (III.1) are
$\Theta_{m=1}^{Brink}$ $=$ $h\frac{\partial}{\partial V}+c\Big({\tanh U}\,\frac{\partial}{\partial X}-(\operatorname{sech}^{2}U)X\frac{\partial}{\partial V}\Big)$ (IV.5)
$+$ $b\Big({\big(U\tanh U-1\big)}\frac{\partial}{\partial X}-\big(\tanh U+U\operatorname{sech}^{2}U\big)X\frac{\partial}{\partial V}\Big)\,.$
$\Theta_{m=2}^{Brink}\!$ $=$ $\!h\frac{\partial}{\partial V}+c\left(\big(3\tanh^{2}U-1\big)\frac{\partial}{\partial X}-\big(6\tanh U\operatorname{sech}^{2}(U)\big)X\frac{\partial}{\partial V}\right)$
$\thinspace{+}b\left(\frac{1}{4}\big(2U-3\tanh U-3U\operatorname{sech}^{2}U\big)\frac{\partial}{\partial X}-\frac{1}{2}\big(1-3\operatorname{sech}^{2}U+3U\operatorname{sech}^{2}U\tanh U\big)X\frac{\partial}{\partial V}\right).\qquad$
V Planar representation of the symmetry generators
The symmetry-generating vectors along the trajectories, shown in FIGs.3 and 4, are conveniently viewed in a co-moving tangent space – which is indeed a plane carried along the trajectory. In FIGs. 58 the trajectory is “hidden” in yellow “blobs” and the coordinates $\xi$ and $\eta$ represent the coefficients of ${\partial}_{X}$ and of ${\partial}_{V}$ of the symmetry generators in (III.1), $\Theta^{Brink}={\bm{\xi}}{\partial}_{X}+\eta{\partial}_{V}$, spelt out in (IV.5)-(IV).
Figure 5: Infinitesimal translations $({\bm{\xi}},\eta)$ along a DM Brinkmann trajectory ${\bm{X}}(U)$ with $\bf m=1$}, shown in the co-moving tangent space. The trajectory is “hidden” in the yellow blob.
FIG.5 shows that the translation vector starts at $\bf U=-\infty$} from $\color[rgb]{.75,.5,.25}{\bm{\xi}}=-1,\eta=0$}. It leaves the caustic point for $\bf U=0$} at ${\bm{X}}=0,\,V=0$ invariant. Arriving into the Afterzone, the translation vectors change sign and end, for $U=+\infty$, at $\color[rgb]{0,1,1}{\bm{\xi}}=+1,{\eta}=0$}. Thus they act again as usual translations – but with reversed sign.
Figure 6: Infinitesimal boosts along along the trajectory “hidden behind the yellow blob” with $\bf m=1$}, shown in the co-moving tangent space.
In FIG.6, the longest arrow (in brown) is both the initial and the final boost vector for $\bf U=\mp\infty$}. The shortest red arrow $({\bm{\xi}}=1,{\eta}=0)$ is reached for $\bf U=0$} at the caustic point ${\bm{X}}=0,V=0$, where the boost shifts all trajectories by the same amount.
Figure 7: Translations act along the $m=2$ DM trajectory by (IV) The trajectory ${\bm{X}}(U)$ with $\bf m=2$} is “hidden” in the yellow blob.
According to FIG.7, the translation generator (with $c=1/2$) is, for $U=$-$\infty$ , $({\bm{\xi}}=1,\,\eta=0)$. FIGs. 3 and 4 confirm that, consistently with (IV.5), the two focal points of the trajectory, at $U_{\mp}$, are left invariant. For $U=+\infty$ the “8-shaped” curve returns to $({\bm{\xi}}=1,\,\eta=0)$, where it started from.
Figure 8: Infinitesimal Carroll boosts in Brinkmann coordinates along a chosen DM trajectory marked by a yellow blob, with wave number $\bf m=2$}.
Boosts along the $m=2$ trajectory are shown in. FIG.8. The longest vector on the left (in brown) is, consistently with (IV.5) and (IV) with $b=1/2$ and $c=2$, the boost for $U=$-$\infty$, and the longest one on the right (also in brown) is for $U=\infty$. The two shortest blue arrows (${\bm{\xi}}=\pm 1,{\eta}=0$) show the boost acting as a translations by $\mp b$ at the caustic points $U_{\mp}$.
VI Conclusion
Our results shed further light on the relation of Carroll symmetry and the Memory Effect : the matrix $P$ with initial conditions (II.16) yields our geodesics. The general expression (III.5) which involves both DM and VM is one of our main results.
When the parameter $m$ is an integer, then the wavezone accommodates $m$ half-waves and we get DM. For $m=2\ell+1$ $P(U)$ is odd and we do get net displacement. For $m=2\ell$, though, $P(U)$ is even and the particle returns, after non-trivial motion in the wave zone, to its initial position : we get DM with no final displacement, consistently with (II.15) and confirmed by FIGs 2, 3a and 4a.
DM trajectories are distinguished by having zero momentum and can be found by putting ${\bm{p}}_{0}=0$ into (III.5). Turning on the momentum, ${\bm{p}}_{0}\neq 0$, distorts the trajectory to VM.
The merit of BJR coordinates is their simplicity Sou73 ; Carroll4GW . For our initial conditions (II.16), the “motion” reduces to a fixed point, given by the conserved boost momentum ${\bm{x}}_{0}={\bf k}$ in (II.10). The entire dynamics is carried by $P(u)$. The price to pay is that the BJR description works only between zeros of the $P$-matrix, and must then be fitted together. These same points, distinguished by the vanishing of $\det(P)$.
The advantage of the Brinkmann form (III.1), which is one of our principal results, is that it replaces the necessarily singular Souriau matrix $S$ Sou73 by the globally defined Sturm-Liouville solution $Q$. As examplified by the Pöschl - Teller profile, it allowed us to extend the locally given Carrollian Killing vectors to entire spacetime. The clue is that the multiplying by $P$ cures the singularity of the Souriau matrix $S$ and thus that of the BJR coordinate system. The Brinkmann trajectories are focused at the zeros of $\det(P)$, as shown in FIG.2.
The Carroll symmetry generators, first identified by Souriau Sou73 , look more complicated in Brinkmann coordinates as their BJR counterparts do. They are however defined globally. The are directly related to geodesics and allow us to unify the Displacement and Velocity Memory effects by using the two independent solutions $P$ and $Q$ of the Sturm-Liouville problem, see (III.5). The intimate relationship of geodesics with Carroll symmetry is highlighted by their common use of the Sturm-Liouville equation GenRelGrav .
Our results go actually beyond $D=1$ dimensional profiles and the Pöschl - Teller example. Similar results hold for the Scarf profile in $D=2$, which yields a good analytic approximation of flyby ZelPol ; Sila-PLB .
The supersymmetric extension MSUSY sheds further light on the Memory Effect. Further details will be presented in a forthcoming comprehensive review ZEHPR .
Acknowledgements.
The authors are indebted to J. Balog for correspondance. PMZ was partially supported by the National Natural Science Foundation of China (Grant No. 11975320). ME is supported by Istanbul Bilgi University research fund (BAP) with grant no: 2024.01.009 and by The Scientific and Technological Research Council of Turkey (TÜBITAK) under grant number 125F021. PMZ and PAH benefitted of TÜBITAK 2221 grant and hospitality at Bilgi University during their visit in Istanbul.

### v2

@@
 Abstract
-The locally defined Carroll symmetry of a gravitational wave is extended to a globally defined one.
+The locally defined Carroll symmetry of a gravitational wave is extended to a globally defined one by switching to Brinkmann coordinates.
 Translations and Carroll boosts associated with two independent globally defined solutions of a Sturm-Liouville equation allow us to describe the motions.
 The Displacement Memory Effect arises for particular choices of the parameters which yield trajectories with zero momentum.
+The relation to the Schwarzian derivative is highlighted.
 We illustrate our general statements by the Pöschl-Teller profile.
@@
 04.20.-q Classical general relativity
-preprint: Global Carroll v1. arXiv: [gr-qc, hep-th]
+preprint: 2510.16762v2 [gr-qc]
 I Introduction
@@
 Another remarkable consequence of (II.15) is that when $\det P(u_{0})=0$ for some $u_{0}$, then all transverse trajectories which start from an $x_{0}\in\mathrm{Ker}\,\big(P(u_{0})\big)$ are focused at the origin, and we get a caustic point BoPi89 ; EZHRev .
-The question will be further discussed in Sec.IV and illustrated in FIG.
+The question will be further discussed in Sec.V and illustrated in FIG.
 2.
@@
 The coefficients of ${\partial}_{X}$ and of ${\partial}_{V}$ are indeed the the symmetry generators $\xi$ and $\eta$, mentioned in the Introduction.
-In Sec.IV they will be spelt out explicitly for Pöschl - Teller .
+In Sec.V they will be spelt out explicitly for Pöschl - Teller .
 The $P$-matrix and its derivative $P^{\prime}$ generate translations, whereas $PS$ and $(PS)^{\prime}$ generate Carroll boosts GenRelGrav .
@@
 is regular for all $U$. Thus (III.1) extends the Carroll symmetry (II.6) naturally from a fixed interval $I_{k}$ to the entire $U$-axis.
-Both $P$ and $Q$ are solutions of the Sturm-Liouville equation (II.3) EZHRev ; Blau24 as it can be verified by a direct calculation.
+The remarkable fact is that both $P$ and $Q$ are solutions of the Sturm-Liouville equation (II.3) EZHRev ; Jibril ; Blau24 , as it can be verified by a direct calculation.
 The Wronskian is $W(P,Q)=P^{\prime}Q-Q^{\prime}P=-1$, therefore $P$ and $Q$ are independent.
@@
 For parameters which correspond to having an integer number of half-waves in the wavezone, we have also $P(+\infty)=\mathop{\rm const}\nolimits$ and we get DM.
-IV Illustration by the Pöschl - Teller profile
+IV Schwarzian aspects
+For simplicity, we work in $D=1$. Our results have an interesting relation with the Schwarzian derivative of a function $f$ of is some real variable $x$ Ovsienko ; Vlahakis ; ZZH ,
+${\mathfrak{S}}\big(f(x)\big)=\frac{f^{\prime\prime\prime}(x)}{f^{\prime}(x)}-\frac{3}{2}\left(\frac{f^{\prime\prime}(x)}{f^{\prime}(x)}\right)^{2}\,.$ (IV.1)
+We note for later use that the Schwarzian depends only on $f^{\prime}$, not on $f$ itself.
+Consider now the Sturm-Liouville equation
+$\varphi^{\prime\prime}(x)-{\mathcal{A}}(x)\varphi=0\,$, (IV.2)
+cf. (II.3) for both $P$ or for $Q$ collectively. Let then $\varphi_{1}(x)$ and $\varphi_{2}(x)$ be two independent solutions and consider their quotient,
+$f(x)=\frac{\varphi_{1}(x)}{\varphi_{2}(x)}\,.$ (IV.3)
+It is then shown in Ovsienko that the Sturm-Liouville profile can be recovered from the solutions in terms of the Schwarzian derivative,
+${{\mathcal{A}}=-{\scriptstyle{\frac{1}{2}}}\,{\mathfrak{S}}(f)\,.}\;\;$ (IV.4)
+Let us now apply these general formulas with the cast $x\leadsto U$, completed by,
+$\varphi_{1}(U)=P(U){\quad\text{\small and}\quad}\varphi_{2}(U)=P(U)S(U)\,.$ (IV.5)
+Thus
+$f=\frac{1}{S}\,.$ (IV.6)
+The Wronskian $W(P,Q)$ is a constant, implying that $f^{\prime}$ in (IV.1) is indeed proportional to $S^{-2}.$ Thus by (IV.4) the profile is the Schwarzian derivative of the inverse of the Souriau matrix,
+${\mathcal{A}}=-\frac{1}{2}\,{\mathfrak{S}}\left(\frac{1}{S}\right)\,.$ (IV.7)
+The form (II.7) of $S$ then implies that $P$ determines also the second solution we called $Q$.
+Remarkably, we can also interchange the cast: choosing
+$\varphi_{1}(U)=Q=P(U)S(U){\quad\text{\small and}\quad}\varphi_{2}(U)=P(U)\,$ (IV.8)
+instead of (IV.5), it is $P$ which drops out, leaving us with the reciprocal relation for
+$\tilde{f}=\frac{Q}{P}=S\,$, (IV.9)
+instead of $f=P/Q=1/S$ as in (IV.6)  [^4] The singularities of $S$ has will be discussed in the next section for PT.
+The new cast (IV.8) then yields, once again, a SL eqn.
+(IV.2) whose profile is now determined by the Souriau matrix  [^5] Eqn.
+(IV.10) can be viewed as a equation for the Souriau matrix $S$.,
+${\mathcal{A}}(U)=-2{\mathfrak{S}}(S)=-2\frac{S^{\prime\prime\prime}(U)}{S^{\prime}(U)}+{3}\left(\frac{S^{\prime\prime}(U)}{S^{\prime(}U)}\right)^{2}\,.$ (IV.10)
+Somewhat surprisingly, inserting here $S$ from (II.7), the previous Sturm-Liouville equation (II.3) with the same profile $\mathcal{A}$} is recovered. These results indicate that the interchange
+$P,\,Q\to Q,\,P\;\;\;\Rightarrow\;\;\;S\to\frac{1}{S}$ (IV.11)
+is a symmetry of the system, as it follows from the reciprocity relation
+${\mathfrak{S}}(f)={\mathfrak{S}}\left(\frac{1}{f}\right)$ (IV.12)
+of the Schwarzian.
+The difference between the two casts comes from the initial conditions.
+Those for $P$ are dictated by the behaviour at $U=$-$\infty$ but those for $Q$ depend also on the integral (II.7).
+V Illustration by the Pöschl - Teller profile
 We illustrate our general theory by the Pöschl - Teller profile PTeller ; Plyushchay09 ; ChaKar ; DM-1 in $D=1$,
-${\mathcal{A}}\equiv{\mathcal{A}}^{PT}(U)=-\dfrac{4m(m+1)}{\cosh^{2}U}\,.$ (IV.1)
+${\mathcal{A}}\equiv{\mathcal{A}}^{PT}(U)=-\dfrac{4m(m+1)}{\cosh^{2}U}\,.$ (V.1)
 The Sturm-Liouville equation (II.3) is solved by Legendre functions. DM trajectories arise when $m$ is an integer, and we get Legendre polynomials, $(-1)^{m}P_{m}(\tanh U)$ DM-1 which have $m$ zeros.
-Turning to BJR by (II.4), the Souriau matrix $S$ in (II.7) is regular between two subsequent zeros of $P$ but diverges at the junction points. For $m=1$, for example,
-$S_{m=1}(u)=u-\coth u\,$ (IV.2)
+Turning to BJR by (II.4), the Souriau matrix $S$ in (II.7) is regular between two subsequent zeros of $P$, but diverges at the junction points. For $m=1$, for example,
+$S_{m=1}(u)=u-\coth u\,$ (V.2)
 obtained for the choice $u^{\pm}_{0}\approx\pm 1.2$, is regular either in $I_{-}<0$ or in $I_{-}>0$, but diverges at $u=0$, as depicted in FIG.1a. Similarly for $m=2$,
-$S_{m=2}=\frac{1}{4}\left(u+\frac{3\sinh\left(2u\right)}{2(2-\cosh(2u))}\right)\,$, (IV.3)
+$S_{m=2}=\frac{1}{4}\left(u+\frac{3\sinh\left(2u\right)}{2(2-\cosh(2u))}\right)\,$, (V.3)
 is singular where the denominator vanishes, as shown in FIG.1b.
-Figure 1: The Sturm-Liouville solutions $\bf P$ in (II.3) are DM trajectories written in Brinkmann coordinates with wave numbers $\bf m=1$} and $\bf m=2$}, respectively.
+Figure 1: For integer $m$ the Sturm-Liouville solutions $\bf P$ in (II.3) are DM trajectories, written in Brinkmann coordinates with wave numbers $\bf m=1$} and $\bf m=2$}, respectively.
 The Souriau matrix S diverges where $\bf P=0$, but the pull-back to Brinkmann $\color[rgb]{0,0.5,0}\bf Q=PS$} is regular for all $U$.
-A second solution of the Sturm-Liouville equation (II.3), $Q_{m}=P_{m}S_{m}$ , is regular for all $U$ but it is a Legendre function with non-DM behavior.
-The solutions $P_{m}$ and $Q_{m}$ are independent when the Wronskian does not vanish, $W(P,Q)\neq 0$.
-For $m=1$ and $2$ we have, for example,
-$\begin{array}[]{llclllc}P_{1}(U)&=&{-}\tanh U\,,&\qquad&P_{2}(U)&=&\frac{1}{2}(3\tanh^{2}U-1)\,,\\[6.0pt] Q_{1}(U)&=&{1-U\tanh U},&\qquad&Q_{2}(U)&=&\frac{1}{2}\big(2U-3\tanh U-3U\operatorname{sech}^{2}U\big)\,.\end{array}$ (IV.4)
+$Q_{m}=P_{m}S_{m}$ , a second solution (III.3) of the Sturm-Liouville equation (II.3) shown in FIG.1, is a Legendre function.
+It is regular for all $U$, vanishes where the Souriau matrix does, and exhibits a VM behaviour where $P_{m}$ is DM.
+The solutions $P_{m}$ and $Q_{m}$ are independent when the Wronskian does not vanish, $W(P,Q)\neq 0$. For $m=1$ and $2$ we have, for example,
+$\begin{array}[]{llclllc}P_{1}(U)&=&{-}\tanh U\,,&\qquad&P_{2}(U)&=&\frac{1}{2}(3\tanh^{2}U-1)\,,\\[6.0pt] Q_{1}(U)&=&{1-U\tanh U},&\qquad&Q_{2}(U)&=&\frac{1}{2}\big(2U-3\tanh U-3U\operatorname{sech}^{2}U\big)\,.\end{array}$ (V.4)
 The two SturmLiouville{Sturm-Liouville\;} solutions are shown in FIG. 1 in red and in green.
@@
 For $m=1$ there is one focal point at the origin, and for $m=2$, we have two of them, determined by $\tanh U_{\pm}=\pm{1}/{\sqrt{3}}$.
-Coming to symmetries of Pöschl - Teller , the coefficients of the Carroll generators $P$ and $Q$ in (III.1) are shown in FIGs. 3 and 4.
+Coming to symmetries, the coefficients of the Carroll generators $P$ and $Q$ in (III.1) are shown in FIGs. 3 and 4.
 (a)                       (b)
@@
 (a)                       (b)
-Figure 4: Infinitesmal Carroll generators (IV) with wave number $\bf m=2$}.
+Figure 4: Infinitesmal Carroll generators (V) with wave number $\bf m=2$}.
 (a) For a translation, the transverse component ${\bf P}_{2}$ is even, and has two zeros, at $U_{\mp}$.
@@
 The globally defined Carroll symmetry generators (III.1) are
-$\Theta_{m=1}^{Brink}$ $=$ $h\frac{\partial}{\partial V}+c\Big({\tanh U}\,\frac{\partial}{\partial X}-(\operatorname{sech}^{2}U)X\frac{\partial}{\partial V}\Big)$ (IV.5)
+$\Theta_{m=1}^{Brink}$ $=$ $h\frac{\partial}{\partial V}+c\Big({\tanh U}\,\frac{\partial}{\partial X}-(\operatorname{sech}^{2}U)X\frac{\partial}{\partial V}\Big)$ (V.5)
 $+$ $b\Big({\big(U\tanh U-1\big)}\frac{\partial}{\partial X}-\big(\tanh U+U\operatorname{sech}^{2}U\big)X\frac{\partial}{\partial V}\Big)\,.$
@@
 $\thinspace{+}b\left(\frac{1}{4}\big(2U-3\tanh U-3U\operatorname{sech}^{2}U\big)\frac{\partial}{\partial X}-\frac{1}{2}\big(1-3\operatorname{sech}^{2}U+3U\operatorname{sech}^{2}U\tanh U\big)X\frac{\partial}{\partial V}\right).\qquad$
-V Planar representation of the symmetry generators
-The symmetry-generating vectors along the trajectories, shown in FIGs.3 and 4, are conveniently viewed in a co-moving tangent space – which is indeed a plane carried along the trajectory.
+VI Representation of the symmetry generators
+The symmetry-generating vectors along the trajectories, shown in FIGs.3 and 4, are conveniently viewed in a co-moving tangent space – which is a plane carried along the trajectory.
 In FIGs.
-58 the trajectory is “hidden” in yellow “blobs” and the coordinates $\xi$ and $\eta$ represent the coefficients of ${\partial}_{X}$ and of ${\partial}_{V}$ of the symmetry generators in (III.1), $\Theta^{Brink}={\bm{\xi}}{\partial}_{X}+\eta{\partial}_{V}$, spelt out in (IV.5)-(IV).
+58 the trajectory is “hidden” in yellow “blobs” and the coordinates $\xi$ and $\eta$ represent the coefficients of ${\partial}_{X}$ and of ${\partial}_{V}$ of the symmetry generators in (III.1), $\Theta^{Brink}={\bm{\xi}}{\partial}_{X}+\eta{\partial}_{V}$, spelt out in (V.5)-(V).
 Figure 5: Infinitesimal translations $({\bm{\xi}},\eta)$ along a DM Brinkmann trajectory ${\bm{X}}(U)$ with $\bf m=1$}, shown in the co-moving tangent space.
@@
 The shortest red arrow $({\bm{\xi}}=1,{\eta}=0)$ is reached for $\bf U=0$} at the caustic point ${\bm{X}}=0,V=0$, where the boost shifts all trajectories by the same amount.
-Figure 7: Translations act along the $m=2$ DM trajectory by (IV) The trajectory ${\bm{X}}(U)$ with $\bf m=2$} is “hidden” in the yellow blob.
+Figure 7: Translations act along the $m=2$ DM trajectory by (V) The trajectory ${\bm{X}}(U)$ with $\bf m=2$} is “hidden” in the yellow blob.
 According to FIG.7, the translation generator (with $c=1/2$) is, for $U=$-$\infty$ , $({\bm{\xi}}=1,\,\eta=0)$.
 FIGs.
-3 and 4 confirm that, consistently with (IV.5), the two focal points of the trajectory, at $U_{\mp}$, are left invariant.
+3 and 4 confirm that, consistently with (V.5), the two focal points of the trajectory, at $U_{\mp}$, are left invariant.
 For $U=+\infty$ the “8-shaped” curve returns to $({\bm{\xi}}=1,\,\eta=0)$, where it started from.
@@
 FIG.8.
-The longest vector on the left (in brown) is, consistently with (IV.5) and (IV) with $b=1/2$ and $c=2$, the boost for $U=$-$\infty$, and the longest one on the right (also in brown) is for $U=\infty$.
+The longest vector on the left (in brown) is, consistently with (V.5) and (V) with $b=1/2$ and $c=2$, the boost for $U=$-$\infty$, and the longest one on the right (also in brown) is for $U=\infty$.
 The two shortest blue arrows (${\bm{\xi}}=\pm 1,{\eta}=0$) show the boost acting as a translations by $\mp b$ at the caustic points $U_{\mp}$.
-VI Conclusion
-Our results shed further light on the relation of Carroll symmetry and the Memory Effect : the matrix $P$ with initial conditions (II.16) yields our geodesics.
+Combining FIGs.
+3 with FIG.5 provides us with the outfolding to 3dimensions, shown in FIG.
+9.
+Figure 9: The translations implemented along the trajectories with $m=1$ and $m=2$.
+The translations implemented along the trajectories vanish where $P(U)=0$. They return to their initial value for $m$ odd and their transverse component changes sign for $m$ even.
+Figure 10: Carroll boosts implemented along the trajectories with $m=1$ and $m=2$.
+Carroll boosts implemented along the trajectories behave at pure translations at P(U)=0P(U)~=~0.
+They return to their initial value for $m$ odd and their transverse component changes sign for $m$ even.
+Conversely, we duly recover the Pöschl - Teller profile (V.1) from the Sturm-Liouville solutions by using the Schwarzian formulas in sect.IV.
+We show this here for the simplest, $m=1$ solution (V.4), with the cast
+$\varphi_{1}(U)=P(U)=-\tanh U,\qquad\varphi_{2}(U)=Q(U)=1-U\tanh U\,.$ (VI.1)
+The Wronskian equals to (1)(-1) everywhere. Then inserting $f^{\prime}(U)=-\cosh^{2}U+1\,$ into the Schwarzian (IV.4) reproduces the $m=1$ profile ${\mathcal{A}}(U)=-\frac{8}{\cosh^{2}(U)}$ in (V.1).
+VII Conclusion
+Our results shed further light on the intimate relation of Carroll symmetry and the Memory Effect : the matrix $P$ with initial conditions (II.16) yields our geodesics.
 The general expression (III.5) which involves both DM and VM is one of our main results.
@@
 These same points, distinguished by the vanishing of $\det(P)$.
-The advantage of the Brinkmann form (III.1), which is one of our principal results, is that it replaces the necessarily singular Souriau matrix $S$ Sou73 by the globally defined Sturm-Liouville solution $Q$.
-As examplified by the Pöschl - Teller profile, it allowed us to extend the locally given Carrollian Killing vectors to entire spacetime.
+The advantage of the Brinkmann form (III.1) is that it replaces the necessarily singular Souriau matrix $S$ Sou73 by the globally defined Sturm-Liouville solution $Q$ – another one of our principal results,
+Examplified by the Pöschl - Teller profile, it allowed us to extend the locally given Carrollian Killing vectors to the whole spacetime.
 The clue is that the multiplying by $P$ cures the singularity of the Souriau matrix $S$ and thus that of the BJR coordinate system.
@@
 They are however defined globally.
-The are directly related to geodesics and allow us to unify the Displacement and Velocity Memory effects by using the two independent solutions $P$ and $Q$ of the Sturm-Liouville problem, see (III.5).
-The intimate relationship of geodesics with Carroll symmetry is highlighted by their common use of the Sturm-Liouville equation GenRelGrav .
-Our results go actually beyond $D=1$ dimensional profiles and the Pöschl - Teller example.
-Similar results hold for the Scarf profile in $D=2$, which yields a good analytic approximation of flyby ZelPol ; Sila-PLB .
-The supersymmetric extension MSUSY sheds further light on the Memory Effect. Further details will be presented in a forthcoming comprehensive review ZEHPR .
+They are also directly related to geodesics and allow us to unify the Displacement and Velocity Memory effects by making use of the two independent solutions $P$ and $Q$ of the Sturm-Liouville problem, see (III.5).
+The intimate relationship of geodesics with Carroll symmetry is highlighted by their common use of the Sturm-Liouville equation EZHRev ; GenRelGrav .
+Figs. 510 illustrate the curious behavior of the Carroll generators along the trajectories.
+Further insight is provided by the Schwarzian framework Ovsienko ; ZZH ; Vlahakis .
+This approach confirms that $Q=PS$ is indeed a global, independent, 2nd solution of the Sturm-Liouville problem which does not involve the locally defined Souriau matrix $S$.
+The supersymmetric extension MSUSY sheds further light on the Memory Effect.
+Our results go beyond $D=1$ dimensional profiles and the Pöschl - Teller example.
+Similar results hold for the Scarf profile in $D=2$, which yields an analytic approximation of flyby ZelPol ; Sila-PLB .
+Details will be presented in a forthcoming comprehensive review ZEHPR .
 Acknowledgements.
 The authors are indebted to J.
-Balog for correspondance.
+Balog, J.
+Ben Achour and Z.
+Silagadze for correspondance.
 PMZ was partially supported by the National Natural Science Foundation of China (Grant No.

### v3

@@
 Abstract
-The locally defined Carroll symmetry of a gravitational wave is extended to a globally defined one by switching to Brinkmann coordinates.
-Translations and Carroll boosts associated with two independent globally defined solutions of a Sturm-Liouville equation allow us to describe the motions.
-The Displacement Memory Effect arises for particular choices of the parameters which yield trajectories with zero momentum.
+The local Carroll symmetry of a gravitational wave found in Baldwin-Jeffery-Rosen coordinates is extended to a globally defined one by switching to Brinkmann coordinates.
+Two independent globally defined solutions of a Sturm-Liouville equation allow us to describe both the symmetries (translations and Carroll boosts) and the geodesic motions.
+One of them satisfies particular initial conditions which imply zero initial momentum, while the other does not.
+Pure displacement arises when the latter is turned off by requiring the momentum to vanish and when the wave parameters take, in addition, some particular values which correspond to having an integer half-wave number.
 The relation to the Schwarzian derivative is highlighted.
@@
 04.20.-q Classical general relativity
-preprint: 2510.16762v2 [gr-qc]
 I Introduction
-Braginsky and Thorne had suggested that gravitational waves might be observed by detecting the displacement of particles initially in rest, called later the Memory Effect BraTho .
-While experimental verification is still in the making, an insight can be gained by taking advantage of the Carroll symmetry Leblond ; SenGupta .
-Baldwin-Jeffery-Rosen (BJR) coordinates BaJe ; Rosen provide a simple description Sou73 ; Carroll4GW , however suffer of being regular only in finite intervals, requiring to glue them together.
-In this paper we show that such complications can be avoided by switching to Brinkmann (B) coordinates Bri by solving a matrix Sturm-Liouville equation, (II.3) below.
-The price to pay is, though, to get more complicated-looking expressions.
-We illustrate our general theory by the Pöschl-Teller profile PTeller ; Plyushchay09 ; ChaKar , which is a very good analytical approximation of the widely studied Gaussian one.
-We follow mostly our review EZHRev which provides the reader also with further references.
-Our notations are: $({\bm{X}},U,V)$ are Brinkmann coordinates and $({\bm{x}},u,v)$ are BJR coordinates on relativistic spacetime.
-$U$ is an affine parameter for geodesics.
-$\Theta^{Brink}={\bm{\xi}}{\partial}_{{\bm{X}}}+\eta{\partial}_{V}$ denote vector fields which generate the infinitesimal isometries on spacetime.
+Braginsky and Thorne BraTho argued that gravitational waves might be observed by detecting the displacement of particles initially in rest, called later the Memory Effect.
+While experimental verification is still in the making, conceptual insight can be gained by taking advantage of the Carroll symmetry LeblondCar ; SenGupta .
+Baldwin-Jeffery-Rosen (BJR) coordinates BaJe ; Rosen provide a simple description Sou73 ; Carroll4GW , however are regular only in finite intervals, requiring to glue together the local results.
+In this paper we show that such complications can be avoided by switching to globally defined Brinkmann (B) coordinates Bri ; DGH91 .
+The transcription requires to solve a matrix Sturm-Liouville equation, (II.6) below.
+The price to pay for globality is to get more complicated-looking expressions, though.
+The advantages and drawbacks of the two coordinate systems is discussed in the Conclusion.
+We illustrate our general theory by the Pöschl-Teller profile PTeller ; Plyushchay09 ; ChaKar , which is a good analytical approximation of the widely studied Gaussian one.
+The book exactsol and our review EZHRev provide the reader with further references.
+Our notations are: $({\bm{X}},U,V)$ are Brinkmann coordinates and $({\bm{x}},u,v)$ are BJR coordinates.
+$U=u$ is an affine parameter for geodesics.
+$\Theta^{Brink}={\bm{\xi}}{\partial}_{{\bm{X}}}+\eta{\partial}_{V}$ denotes collectively the vector fields which generate the infinitesimal isometries of spacetime.
+Its explicit form will be spelt out in (III.1).
 II From Brinkmann to BJR and back
-Plane gravitational waves are conveniently described by using globally defined Brinkmann coordinates Bri ; exactsol in terms of which the metric is,
+Plane gravitational waves are conveniently described by using globally defined Brinkmann coordinates Bri ; DGH91 ; exactsol in terms of which the metric is,
 $g_{\mu\nu}dX^{\mu}dX^{\nu}=d{\bm{X}}^{2}+2dUdV+K_{ij}X^{i}X^{j}\,dU^{2}\,$, (II.1)
-where ${\bm{X}}=(X^{1},\dots,X^{D})$ parametrizes the transverse plane, endowed with the flat Euclidean metric $d{\bm{X}}^{2}=\delta_{ij}\,dX^{i}dX^{j}$.
-$U$ and $V$ are light-cone coordinates and $(K_{ij})$ we call the profile is a symmetric $D\times D$ matrix whose entries depend only on $U$.
-Henceforth we focus our attention at $D=1$ or $D=2$ and assume, for simplicity, that is diagonal.
-In $D=2$ we take
+where the ${\bm{X}}=(X^{1},\dots,X^{D})$ are coordinates of the transverse plane with the flat Euclidean metric $d{\bm{X}}^{2}=\delta_{ij}\,dX^{i}dX^{j}$.
+$U$ and $V$ are light-cone coordinates related to the usual relativistic coordinates according to $U=\frac{1}{\sqrt{2}}(Z-T)$ and $V=\frac{1}{\sqrt{2}}(Z+T)$.
+The profile $(K_{ij})$ is a symmetric $D\times D$ matrix whose entries depend only on $U$.
+Henceforth we focus our attention at $D=1$ or $D=2$ and assume, for simplicity, that the wave is linearly polarized so that $(K_{ij})$ is diagonal.
+Thus in $D=2$, which is our main interest in this paper,
 $K_{ij}X^{i}X^{j}={\mathcal{A}}(U)\Big((X^{+})^{2}-(X^{-})^{2}\Big)\,$, (II.2)
-which yields a vacuum solution of the Einstein equations Bri ; exactsol .
-We consider “sandwich waves” BoPi89 , whose profile ${\mathcal{A}}(U)$ vanishes outside an interval $U_{i}<U<U_{f}$.
-The $V$-equation follows from the transverse ones EZHRev and will henceforth not be studied.
-After the wave has passed, particles initially at rest exhibit the Velocity Memory Effect (VM) : they move with constant velocity.
-Under certain “quantization” conditions the outgoing velocity can vanish, though, and we get the Displacement Memory Effect (DM) ZelPol ; DM-1 ; DM-2 ; Sila-PLB , which might well play a rôle in future observations BraTho ; DM-2 .
+which yields a vacuum solution of the Einstein equations for any ${\mathcal{A}}(U)$ Bri ; exactsol ; DGH91 ; EZHRev .
+We consider only “sandwich waves” BoPi89 , whose profile ${\mathcal{A}}(U)$ vanishes outside an interval $U_{i}<U<U_{f}$.
+Choosing $U$ as an affine parameter, the equations of motion of the transverse coordinates are of the Sturm-Liouville form,
+$\dfrac{d^{2}{\bm{X}}}{dU^{2}}-\left(\begin{array}[]{lr}{\mathcal{A}}&0\\ 0&-{\mathcal{A}}\end{array}\right)\,{\bm{X}}=0\,.$ (II.3)
+The geodesic equations have an additional constant of the motion called the Jacobi invariant exactsol ,
+$e=\frac{1}{2}\,g_{\mu\nu}\,\frac{d{X}^{\mu}}{dU}\frac{d{X}^{\nu}}{dU}\,.$ (II.4)
+The transverse equations (II.3) do not depend on $e$. In this paper we limit our attention at the lightlike case $e=0$ for which the “vertical” coordinate $V$ satisfies
+$\dfrac{d^{2}V}{dU^{2}}+\dfrac{1}{2}\displaystyle{\frac{d{\mathcal{A}}}{dU\,}}\Big((X^{1})^{2}-(X^{2})^{2}\Big)+2{\mathcal{A}}\Big(X^{1}\dfrac{dX^{1}}{dU\,}-X^{2}\dfrac{dX^{2}}{dU\,}\Big)=0\,.$ (II.5)
+It corresponds to the horizontal lift of (II.3) and can be solved once the solutions of the latter are known.
+Henceforth we focus our attention at (II.3).
+The timelike (massive) case, $e<0$, is discussed, e.g.
+in EDAHK ; DM-1 .
+The equations above constitute a model for particle dynamics in a gravitational wave including flyby ZelPol .
+After the wave has passed, particles initially at rest exhibit the Velocity Memory Effect (VM) Ehlers ; AiBalasin ; EZHRev  : outside the Wavezone, they move with constant velocity.
+Under certain “quantization” conditions the outgoing velocity can vanish, though, and we get the Displacement Memory Effect (DM) ZelPol ; DM-1 ; DM-2 ; Sila-PLB .
 Plane gravitational waves have long been known to have a $2D+1$ parameter symmetry group composed of $D+1$ translations, completed by $D$ (rather mysterious) transformations BoPiRo ; Sou73 .
-More recently Carroll4GW , this group was identified with the Carroll group Leblond ; SenGupta , as young Lévy-Leblond called it jokingly.
-Its algebraic structure is readily determined by switching to Baldwin et al (BJR) coordinates BaJe ; Rosen ; Sou73 which are however only local.
-This note sheds further light at the Carroll – Memory relation using globally defined Brinkmann coordinates.
+More recently Carroll4GW , this group was identified with the Carroll group LeblondCar ; SenGupta , as young Lévy-Leblond called it jokingly.
+Its algebraic structure is readily determined by switching to Baldwin-Jeffery-Rosen (BJR) coordinates BaJe ; Rosen ; Sou73 which are however only local.
+This note sheds further light at the Carroll — Memory relation using globally defined Brinkmann coordinates.
 We start with considering real solutions of the Sturm-Liouville equations
-${P}^{\prime\prime}={\mathcal{A}}(U)P,\qquad(P^{T}){P}^{\prime}=(P^{T})^{\prime}P\,$ (II.3)
-for a $D{\times}D$ matrix $P(U)=\big(P_{ij}(U)\big)$ EZHRev . The BJR coordinates $({\bm{x}},u,v)$, are then obtained by
-${{\bm{X}}}=P(u){\bm{x}},\qquad U=u,\qquad V=v-\frac{1}{4}{\bm{x}}\cdot\big(P^{T}{}P\big)^{\prime}{\bm{x}}\,$, (II.4)
-in terms of which the metric is,
-$g_{\mu\nu}dx^{\mu}dx^{\nu}=\big(P(u)^{T}P(u)\big)_{ij}(u)\,dx^{i}dx^{j}+2du\,dv\,.$ (II.5)
-Carroll symmetry consists, in BJR coordinates, of “vertical” ($v$) and transverse ($\bm{x}$}) translations with parameter $h$ and $\bf c$}, completed by Carroll boosts with parameter $\bf b$} Sou73 ; Carroll4GW ,
-${\bm{x}}\to{\bm{x}}+{\bf c}+S(u)\,{\bf b}\,,\qquad u\to u,\qquad v\to v-{\bf b}\cdot{\bm{x}}-{{\scriptstyle{\frac{1}{2}}}}{\bf b}\cdot{}S(u)\,{\bf b}+h\,$, (II.6)
-where the symmetric $2\times 2$ matrix
-$S(u)=\int^{u}_{u_{0}}\!\left(P^{T}P\right)^{-1}\!(t)dt$ (II.7)
-is referred to as the Souriau matrix Sou73 ; Carroll4GW ; EZHRev . Infinitesimal boosts parametrized with $b_{j}$ are,
-$\theta^{BJR}=h\partial_{v}+c^{j}\partial_{j}+b_{j}(S^{ji}\partial_{i}-x^{j}\partial_{v})\,.$ (II.8)
-Finding the trajectories is then straightforward. Noether’s theorem provides us with $2D+1$ associated conserved quantities Sou73 ; Carroll4GW ,
-${\bm{p}}={\bm{p}}_{0}=(P^{T}P)(u)\,{{\bm{x}}}^{\prime}(u),\qquad{\bf k}={\bf k}_{0}={\bm{x}}(u)-S(u)\,{\bm{p}}\,$, (II.9)
-interpreted as conserved linear and boost-momentum, supplemented by $m={u}^{\prime}=1$. Then the conserved quantities determine the geodesic,
-${\bm{x}}(u)={\bf k}_{0}+S(u)\,{\bm{p}}_{0}\,.$ (II.10)
-The subtlety comes from that the BJR coordinates are defined only in intervals $I_{k}=[u_{k-1},u_{k}]$ between the (mandatory Sou73 ; Carroll4GW ) zeros of the determinant,
-$\det P(u_{k})=0\,$, (II.11)
-which play a fundamentally important rôle in our investigation, as we shall see.
-The BJR coordinates are thus only local : both the metric (II.5) and the Souriau matrix $S$ in (II.7) are singular at the junction points $u_{k}$, as will be illustrated in FIG.1 for the Pöschl - Teller profile.
-The singularity problem will be resolved in sect.
-III by switching to Brinkmann coordinates.
-Now we set $P^{T}=P$ and we consider $D=1$ (when $P$ is a scalar function) for simplicity, although our results and conclusions could be extended to non-diagonal profiles and to any dimension.
-Let us assume that $u<u_{1}$, the first zero of $\det(P)$.
-Requiring that our particles be in rest before the wave arrives,
-${{\bm{x}}}(-\infty)=x_{0}{\quad\text{\small and}\quad}{{\bm{x}}}^{\prime}(-\infty)=0\,$, (II.12)
-imply that the momentum vanishes in $I_{1}$,
-${\bm{p}}=0\,.$ (II.13)
-The Souriau term is thus switched off from (II.10) and the transverse BJR “trajectory” is merely a fixed point.
-${\bm{x}}(u)={\bm{x}}_{0}=\mathop{\rm const}\nolimits\,$, (II.14)
-confirming, in BJR coordinates, the “no motion for Carroll” maxim Leblond . In contrast, the Brinkmann trajectory is non-trivial : (II.4) yields,
-${\bm{X}}(U)=P(u)\,{\bm{x}}_{0}\,$, (II.15)
-with $U=u$. The initial conditions (II.12) then require :
-$P(-\infty)=1{\quad\text{\small and}\quad}P^{\prime}(-\infty)=0\,.$ (II.16)
-The Brinkmann trajectory, given by the $P$-matrix, may look quite complicated, depending on the Sturm-Liouville solution $P$ of (II.4) EZHRev .
-For appropriate “quantized” values of the wave parameter the wave zone contains an integer number of half-waves; then $P(u)$ tends to a constant matrix also for $u\to\infty$ : DM is obtained DM-1 ; DM-2 ; Sila-PLB .
-We emphasise that the investigations above are a priori valid only before the first zero of $\det{P}$ and should then be restarted until the next zero, and so on.
-Another remarkable consequence of (II.15) is that when $\det P(u_{0})=0$ for some $u_{0}$, then all transverse trajectories which start from an $x_{0}\in\mathrm{Ker}\,\big(P(u_{0})\big)$ are focused at the origin, and we get a caustic point BoPi89 ; EZHRev .
-The question will be further discussed in Sec.V and illustrated in FIG.
-2.
+${P}^{\prime\prime}={\mathcal{A}}(U)P,\qquad(P^{T}){\,}{P}^{\prime}=(P^{T})^{\prime}{\,}P\,$ (II.6)
+for a $D{\times}D$ matrix $P(U)=\big(P_{ij}(U)\big)$ EZHRev . The prime here denotes $d/du\equiv d/dU$. BJR coordinates $({\bm{x}},u,v)$ are then obtained by
+${{\bm{X}}}=P(u){\bm{x}},\qquad U=u,\qquad V=v-\frac{1}{4}{\bm{x}}\cdot\big(P^{T}{}P\big)^{\prime}{\bm{x}}\,$, (II.7)
+in terms of which the metric takes the BJR form,
+$g_{\mu\nu}dx^{\mu}dx^{\nu}=\left(P(u)^{T}P(u)\right)_{ij}(u)\,dx^{i}dx^{j}+2du\,dv\,.$ (II.8)
+Carroll symmetry written in BJR coordinates is composed of “vertical” ($v$) and transverse ($\bm{x}$}) translations with parameters $h$ and $\bf c$}, respectively, completed by Carroll boosts with parameter $\bf b$} Sou73 ; Carroll4GW ,
+${\bm{x}}\to{\bm{x}}+{\bf c}+S(u)\,{\bf b}\,,\qquad u\to u,\qquad v\to v-{\bf b}\cdot{\bm{x}}-{{\scriptstyle{\frac{1}{2}}}}{\bf b}\cdot{}S(u)\,{\bf b}+h\,.$ (II.9)
+The symmetric $2\times 2$ matrix
+$S(u)=\int^{u}_{u_{0}}\!\left(P^{T}P\right)^{-1}\!(t)dt$ (II.10)
+here is referred to as the Souriau matrix Sou73 ; Carroll4GW ; EZHRev .
+$S$ is related to but is slightly different from the time redefinition considered, in the isotropic case, by Arnold Arnold ; ZurabNote ; Aniso ; ZZH .
+Details will be as discussed elsewhere ZEHPR .
+The infinitesimal symmetries are, with some abuse of notation,
+$\theta^{BJR}=h\partial_{v}+c^{j}\partial_{j}+b_{j}\big(S^{ji}\partial_{i}-x^{j}\partial_{v}\big)\,.$ (II.11)
+The covariantly constant Killing vector ${\partial}_{v}$ here generates “vertical” translations – a hallmark of Brinkmann metrics Bri ; DGH91 .
+The $c_{i}$ generate translations and the $b_{j}$ generate Carroll boosts.
+For $u=u_{0}$, the integral (II.10) vanishes and we recover the usual Carroll action in LeblondCar .
+Finding the geodesic trajectories is then straightforward by using the symmetry. Noether’s theorem provides us with associated conserved quantities Sou73 ; Carroll4GW ,
+${\bm{p}}={\bm{p}}_{0}=(P^{T}P)(u)\,{{\bm{x}}}^{\prime}(u),\qquad{\bf k}={\bf k}_{0}={\bm{x}}(u)-S(u)\,{\bm{p}}\,$, (II.12)
+interpreted as conserved linear and boost-momentum, respectively, supplemented by the non-relativistic mass that we scale to $1$.
+It is generated by the Killing vector ${\partial}_{v}$.
+The conserved quantities determine the geodesic trajctories  [^4] A non-vanishing Jacobi invariant $e\neq 0$ would merely add a linear term, $eu$ to the r.h.s., EDAHK ; DM-1 .,
+${\bm{x}}(u)={\bf k}_{0}+S(u)\,{\bm{p}}_{0}\,,\quad v(u)=-\frac{1}{2}{\bm{x}}.{\bm{p}}+v_{0}\,.$ (II.13)
+So far so good.
+However a subtlety comes from that the BJR coordinates are defined only in intervals $I_{k}=[u_{k-1},u_{k}]$ between mandatory Sou73 ; Carroll4GW zeros of the determinant [^5] In $D=1$, $P(U)$ is just a scalar function.,
+$\det P(u_{k})=0\,.$ (II.14)
+The matrix $P$ plays a fundamentally important rôle in our investigation, as we shall see.
+Both the metric (II.8) and the Souriau matrix $S$ (II.10) are singular at the junction points $u_{k}$, as is manifest in FIG.2 for the Pöschl - Teller profile.
+The singularity will be removed by switching to Brinkmann coordinates.
+Although our results could be extended to non-diagonal profiles and to any dimension, henceforth we focus our attention at the diagonal case, $P^{T}=P$, and to D=2D=2i.e.,  to the physically relevant $4$-dimensional spacetime.
+Occasionally, we consider also toy examples in $D=1$, though.
+Let us assume that $u<u_{1}$, the first zero of $\det(P)$. Requiring that our particles be in rest before the wave arrives,
+${{\bm{x}}}(-\infty)=x_{0}{\quad\text{\small and}\quad}{{\bm{x}}}^{\prime}(-\infty)=0\,$, (II.15)
+implies that the momentum vanishes in $I_{1}$,
+${\bm{p}}=0\,.$ (II.16)
+The Souriau term is thus switched off from (II.13) and the transverse BJR “trajectory” is, for vanishing Jacobi invariant, $e=0$ in (II.4), merely a fixed point  [^6] A non-vanishing Jacobi in $e\neq 0$ would merely shift $v(u)$ by a linear-in-$u$ term EDAHK ; DM-1 .,
+${\bm{x}}(u)={\bm{x}}_{0}=\mathop{\rm const}\nolimits\,,\qquad v(u)=v_{0}=\mathop{\rm const}\nolimits\,$, (II.17)
+consistently with (II.13).
+Eqn (II.17) confirms the “no motion for Carroll” wisdom LeblondCar   [^7] Motion is possible for multi-particle systems MultiCarroll , and for planar models with “exotic” double central extension when the particles follow the Hall law MARSOT ; ExoPeierls ..
+However, unlike as for BJR, the (lightlike) Brinkmann trajectory is non-trivial : (II.7) yields,
+${\bm{X}}(U)=P(u)\,{\bm{x}}_{0}\,,\qquad V(U)=-\frac{1}{2}{\bm{X}}.{\bm{X}}^{\prime}+V_{0}\,$, (II.18)
+with $U=u$ and $V_{0}=v_{0}$. The initial conditions (II.15) then require :
+$P(-\infty)={\rm Id}{\quad\text{\small and}\quad}P^{\prime}(-\infty)=0\,.$ (II.19)
+The Brinkmann trajectory, given by the action of the $P$-matrix, (II.18), may look indeed quite complicated, depending on the Sturm-Liouville solution $P$ of (II.6) EZHRev .
+For randomly chosen wave parameters we get VM Ehlers ; AiBalasin ; EZHRev  : the particles fly apart with constant velocity.
+For appropriately “quantized” values of the wave parameter, though, the wave zone contains an integer number of half-waves; $P(u)$ then tends to a constant matrix also for $u=U\to\infty$ and DM is obtained DM-1 ; DM-2 ; Sila-PLB .
+We emphasise, though, that the investigations above are a priori valid only before the first zero of $\det{P}$ and should then be restarted until the next zero, and so on, as it will be seen in sec.V for Pöschl - Teller .
+Another remarkable consequence of (II.18) is that when $\det P(u_{0})=0$ for some $u_{0}$, then all transverse trajectories which start from $x_{0}\in\mathrm{Ker}\,\big(P(u_{0})\big)$ are focused at the origin : we get a caustic point BoPi89 ; EZHRev , as it will be illustrated in FIG.
+3.
 III Carroll symmetry in Brinkmann coordinates
-Simple as they are, BJR coordinates (II.4) suffer from being valid only in intervals between subsequent zeros of $P$, and the adjacent expressions should be glued together – which may be laborious.
+Simple as they are, BJR coordinates (II.7) suffer from being valid only in intervals between subsequent zeros of $det(P)$.
+The adjacent expressions should be glued together – which may be laborious.
 A global approach can conveniently be given instead by using Brinkmann coordinates.
 For simplicity we restrict our attention at $D=1$ transverse dimension and start with a fixed interval $I_{k}$.
-A straightforward computation then shows that pulling back the BJR expression (II.6) by using (II.4) [or its infinitesimal form (II.8)] yields the Carroll generators written in Brinkmann coordinates,
+A straightforward computation then shows that pulling back the BJR expression (II.9) by (II.7) [or its infinitesimal form (II.11)] yields the Carroll generators written in Brinkmann coordinates,
 ${\Theta^{Brink}=h\frac{\partial}{\partial V}+c\left(P\frac{\partial}{\partial X}-P^{\prime}X\frac{\partial}{\partial V}\right)+b\left((PS)\frac{\partial}{\partial X}-(PS)^{\prime}X\frac{\partial}{\partial V}\right)\,,}\;\;$ (III.1)
-The coefficients of ${\partial}_{X}$ and of ${\partial}_{V}$ are indeed the the symmetry generators $\xi$ and $\eta$, mentioned in the Introduction.
-In Sec.V they will be spelt out explicitly for Pöschl - Teller .
-The $P$-matrix and its derivative $P^{\prime}$ generate translations, whereas $PS$ and $(PS)^{\prime}$ generate Carroll boosts GenRelGrav .
-Their only nonzero commutator yields a vertical translation Blau24 ,
+which could also be verified independently, by checking the equation $L_{Y}g=0$ Torre ; GenRelGrav .
+The coefficients of ${\partial}_{X}$ and of ${\partial}_{V}$ are indeed the symmetry generators $\xi$ and $\eta$ mentioned in the Introduction.
+They will be spelt out explicitly for Pöschl - Teller in Sec.V.
+The $P$-matrix and its derivative $P^{\prime}$ generate translations, whereas $Q=PS$ and $Q^{\prime}=(PS)^{\prime}$ generate Carroll boosts GenRelGrav .
+${\partial}_{V}$ generates the non-relativistic mass scaled to $1$, as does for BJR.
+The only nonzero commutator yields a vertical translation Blau24 ,
 $\Big[\underbrace{P\frac{\partial}{\partial X}-P^{\prime}X\frac{\partial}{\partial V}}_{translation}\ ,\ \underbrace{PS\frac{\partial}{\partial X}-(PS)^{\prime}X\frac{\partial}{\partial V}}_{boost}\Big]=-\partial_{V}\,.$ (III.2)
 The symmetry is thus the Heisenberg algebra.
-$P$ thus plays a double role : by (II.15) it determines the B trajectory, whereas (III.1) says that it also generates translations.
-The B trajectory is obtained indeed from $x_{0}=\mathop{\rm const}\nolimits$ by a (generally $U$-dependent) translation.
-The Brinkmann form (III.1) has an important advantage w.r.t. the BJR expression in (II.6) : multiplication by $P$ supresses the singularity of Souriau matrix $S$ Sou73 ; EZHRev  :
+$P(U)$ thus plays a double role : for DM parameters it determines the complicated-looking Brinkmann trajectory in (II.18), whereas by (III.1) it also generates translations.
+For non-DM parameters, the situation is even more complicated, as will be seen in (III.6).
+The Brinkmann form (III.1) has an important advantage w.r.t. the BJR expression (II.9) : left multiplication by $P$ removes the singularity of the Souriau matrix $S$ Sou73 ; EZHRev  :
 Q(U)=P(U)S(U)Q(U)=P(U)S(U)\, (III.3)
-is regular for all $U$. Thus (III.1) extends the Carroll symmetry (II.6) naturally from a fixed interval $I_{k}$ to the entire $U$-axis.
-The remarkable fact is that both $P$ and $Q$ are solutions of the Sturm-Liouville equation (II.3) EZHRev ; Jibril ; Blau24 , as it can be verified by a direct calculation.
-The Wronskian is $W(P,Q)=P^{\prime}Q-Q^{\prime}P=-1$, therefore $P$ and $Q$ are independent.
-The expression (III.1) could also be verified independently, by checking the symmetry equation $L_{Y}g=0$ Torre ; GenRelGrav .
-Contracting the Brinkmann metric (II.1) with (III.1) provides us with,
-${\bm{p}}_{0}=P{\bm{X}}^{\prime}-\big(P\big)^{\prime}{\bm{X}}{\quad\text{\small and}\quad}{\bf k}_{0}=-Q{\bm{X}}^{\prime}+\big(Q\big)^{\prime}{\bm{X}}\,$, (III.4)
+is regular for all $U$.
+In conclusion, switching form BJR to Brinkmann, (III.1),extends the Carroll symmetry (II.9) from an interval $I_{k}$ to the entire $U$-axis.
+A remarkable observation says that both $P$ and $Q$ are solutions of the Sturm-Liouville equation (II.6),
+${Q}^{\prime\prime}={\mathcal{A}}(U)Q,\qquad(Q^{T}){\,}{Q}^{\prime}=(Q^{T})^{\prime}{\,}Q\,$, (III.4)
+as it can also be verified by a direct calculation.
+$P$ and $Q$ have different initial conditions, though EZHRev ; Jibril ; Blau24 .
+The two solutions are independent when their Wronskian, $W(P,Q)=P^{\prime}Q-Q^{\prime}P$, does not vanish.
+Contracting the Brinkmann metric (II.1) with the Carroll generators (III.1) provides us with,
+${\bm{p}}_{0}=P{\bm{X}}^{\prime}-\big(P\big)^{\prime}{\bm{X}}{\quad\text{\small and}\quad}{\bf k}_{0}=-Q{\bm{X}}^{\prime}+\big(Q\big)^{\prime}{\bm{X}}\,$, (III.5)
 where ${\bm{p}}_{0}$ and ${\bf k}_{0}$ are, a priori, new constants of the motion.
-However expressing them in BJR terms, (II.4), the previous expressions (II.9) are recovered; as anticipated by our notation.
-The Sturm-Liouville solutions $P$ and $Q$ play also yet another rôle, though : either combining (II.10) with (II.4) or solving (III.4) directly yields the general Brinkmann trajectory,
-${{\bm{X}}(U)=P(U)\,{\bf k}_{0}+Q\,{\bm{p}}_{0}\,.}\;\;$ (III.5)
-which is our second main result.
-When ${\bm{p}}_{0}=0$ cf.
-(II.13), then the $Q$ - term is switched off and we recover (II.15) with ${\bm{x}}_{0}={\bf k}_{0}$.
-For parameters which correspond to having an integer number of half-waves in the wavezone, we have also $P(+\infty)=\mathop{\rm const}\nolimits$ and we get DM.
+However expressing them in BJR terms using (II.7), the previous expressions (II.12) are recovered, as anticipated by our notation.
+The Sturm-Liouville solutions $P$ and $Q$ play yet another rôle, though : either combining (II.13) with (II.7) or solving (III.5) directly yields the general Brinkmann trajectory,
+${{\bm{X}}(U)=P(U)\,{\bf k}_{0}+Q(U)\,{\bm{p}}_{0}\,.}\;\;$ (III.6)
+Remember that $P$ satisfies the initial conditions (II.19).
+For vanishing momentum ${\bm{p}}_{0}=0$, cf.
+(II.16), the VM term $Q$ is switched off and we recover (II.18) with ${\bm{x}}_{0}={\bf k}_{0}$.
+The corresponding Brinkmann trajectory may or may not be DM yet, though, depending on the wave profile.
+For parameters for which we have an integer number of half-waves in the wavezone, we do get DM : $P(+\infty)=\mathop{\rm const}\nolimits$, though DM-1 ; DM-2 ; Sila-PLB .
+The case of nonzero initial momentum, ${\bm{p}}_{0}\neq 0$, will be studied in ZEHPR .
 IV Schwarzian aspects
-For simplicity, we work in $D=1$. Our results have an interesting relation with the Schwarzian derivative of a function $f$ of is some real variable $x$ Ovsienko ; Vlahakis ; ZZH ,
+For simplicity, we work in $D=1$. Our results have an interesting relation with the Schwarzian derivative defined for a function $f$ as Ovsienko ; Vlahakis ; ZZH ,
 ${\mathfrak{S}}\big(f(x)\big)=\frac{f^{\prime\prime\prime}(x)}{f^{\prime}(x)}-\frac{3}{2}\left(\frac{f^{\prime\prime}(x)}{f^{\prime}(x)}\right)^{2}\,.$ (IV.1)
@@
 $\varphi^{\prime\prime}(x)-{\mathcal{A}}(x)\varphi=0\,$, (IV.2)
-cf. (II.3) for both $P$ or for $Q$ collectively. Let then $\varphi_{1}(x)$ and $\varphi_{2}(x)$ be two independent solutions and consider their quotient,
+cf. (II.6) for both $P$ or for $Q$ collectively. Let then $\varphi_{1}(x)$ and $\varphi_{2}(x)$ be two independent solutions and consider their quotient,
 $f(x)=\frac{\varphi_{1}(x)}{\varphi_{2}(x)}\,.$ (IV.3)
@@
 ${\mathcal{A}}=-\frac{1}{2}\,{\mathfrak{S}}\left(\frac{1}{S}\right)\,.$ (IV.7)
-The form (II.7) of $S$ then implies that $P$ determines also the second solution we called $Q$.
+The form (II.10) of $S$ then implies that $P$ determines also the second solution we called $Q$.
 Remarkably, we can also interchange the cast: choosing
@@
 $\tilde{f}=\frac{Q}{P}=S\,$, (IV.9)
-instead of $f=P/Q=1/S$ as in (IV.6)  [^4] The singularities of $S$ has will be discussed in the next section for PT.
-The new cast (IV.8) then yields, once again, a SL eqn.
-(IV.2) whose profile is now determined by the Souriau matrix  [^5] Eqn.
-(IV.10) can be viewed as a equation for the Souriau matrix $S$.,
+instead of $f=P/Q=1/S$ as in (IV.6).
+The new cast (IV.8) then yields, once again, a SL eqn. (IV.2) whose profile is determined by the Souriau matrix,
 ${\mathcal{A}}(U)=-2{\mathfrak{S}}(S)=-2\frac{S^{\prime\prime\prime}(U)}{S^{\prime}(U)}+{3}\left(\frac{S^{\prime\prime}(U)}{S^{\prime(}U)}\right)^{2}\,.$ (IV.10)
-Somewhat surprisingly, inserting here $S$ from (II.7), the previous Sturm-Liouville equation (II.3) with the same profile $\mathcal{A}$} is recovered. These results indicate that the interchange
+which can indeed be viewed as an equation for $S$.
+Somewhat surprisingly, inserting here $S$ from (II.10), the previous Sturm-Liouville equation (II.6) with the same profile $\mathcal{A}$} is recovered.
+These results indicate that the interchange
 $P,\,Q\to Q,\,P\;\;\;\Rightarrow\;\;\;S\to\frac{1}{S}$ (IV.11)
@@
 ${\mathfrak{S}}(f)={\mathfrak{S}}\left(\frac{1}{f}\right)$ (IV.12)
-of the Schwarzian.
-The difference between the two casts comes from the initial conditions.
-Those for $P$ are dictated by the behaviour at $U=$-$\infty$ but those for $Q$ depend also on the integral (II.7).
+of the Schwarzian. The difference between the two casts comes from the initial conditions. Those for $P$ are dictated by the behaviour at $U=$-$\infty$ but those for $Q$ involve also $S$ in (II.10).
 V Illustration by the Pöschl - Teller profile
 We illustrate our general theory by the Pöschl - Teller profile PTeller ; Plyushchay09 ; ChaKar ; DM-1 in $D=1$,
-${\mathcal{A}}\equiv{\mathcal{A}}^{PT}(U)=-\dfrac{4m(m+1)}{\cosh^{2}U}\,.$ (V.1)
-The Sturm-Liouville equation (II.3) is solved by Legendre functions. DM trajectories arise when $m$ is an integer, and we get Legendre polynomials, $(-1)^{m}P_{m}(\tanh U)$ DM-1 which have $m$ zeros.
-Turning to BJR by (II.4), the Souriau matrix $S$ in (II.7) is regular between two subsequent zeros of $P$, but diverges at the junction points. For $m=1$, for example,
+${\mathcal{A}}\equiv{\mathcal{A}}^{PT}(U)=-\dfrac{m(m+1)}{\cosh^{2}U}\,.$ (V.1)
+The Sturm-Liouville equation (II.6) is solved by Legendre functions. DM trajectories are obtained when $m$ is an integer and we get Legendre polynomials,
+$P(U)=(-1)^{m}P_{m}(\tanh U)\,.$ (V.2)
+The general geodesic, (III.6), combines the DM solution $P(U)$ and the non-polynomial $Q(U)$ in (III.3), shown FIG.1.
+Figure 1: The solutions of the Sturm-Liouville eqn.
+(II.6) for Pöschl - Teller , shown for DM parameters.
+With initial conditions (II.19) we get Legendre polynomials, $P(U)=P_{m}(\tanh U)$.
+The second, independent solution, $Q(U)=Q_{m}(\tanh U)$, is a Legendre function which does not satisfy the initial condition (II.19).
+Turning to BJR cf. (II.7), the Souriau matrix $S$ in (II.10) is regular between two subsequent zeros of $P$ but diverges at the junction points. For $m=1$, for example,
 $S_{m=1}(u)=u-\coth u\,$ (V.3)
-obtained for the choice $u^{\pm}_{0}\approx\pm 1.2$, is regular either in $I_{-}<0$ or in $I_{-}>0$, but diverges at $u=0$, as depicted in FIG.1a. Similarly for $m=2$,
-$S_{m=2}=\frac{1}{4}\left(u+\frac{3\sinh\left(2u\right)}{2(2-\cosh(2u))}\right)\,$, (V.3)
-is singular where the denominator vanishes, as shown in FIG.1b.
-Figure 1: For integer $m$ the Sturm-Liouville solutions $\bf P$ in (II.3) are DM trajectories, written in Brinkmann coordinates with wave numbers $\bf m=1$} and $\bf m=2$}, respectively.
+obtained for the choice $u^{\pm}_{0}\approx\pm 1.2$, which is regular both in $I_{-}=\big\{u<0\big)\}$ and in $I_{+}=\big\{u>0\big)\}$, but diverges at $u=0$, as depicted in FIG.2a.
+Similarly for $m=2$ shown in FIG.2b,
+$S_{m=2}=\frac{1}{4}\left(u+\frac{3\sinh\left(2u\right)}{2(2-\cosh(2u))}\right)\,$ (V.4)
+is singular where the denominator vanishes.
+Figure 2: The Sturm-Liouville solutions $\bf P$ of (II.6) shown with wave numbers $\bf m=1$} and $\bf m=2$} for Pöschl - Teller in $D=1$.
+They yield also DM trajectories written in Brinkmann coordinates.
 The Souriau matrix S diverges where $\bf P=0$, but the pull-back to Brinkmann $\color[rgb]{0,0.5,0}\bf Q=PS$} is regular for all $U$.
-$Q_{m}=P_{m}S_{m}$ , a second solution (III.3) of the Sturm-Liouville equation (II.3) shown in FIG.1, is a Legendre function.
-It is regular for all $U$, vanishes where the Souriau matrix does, and exhibits a VM behaviour where $P_{m}$ is DM.
-The solutions $P_{m}$ and $Q_{m}$ are independent when the Wronskian does not vanish, $W(P,Q)\neq 0$. For $m=1$ and $2$ we have, for example,
-$\begin{array}[]{llclllc}P_{1}(U)&=&{-}\tanh U\,,&\qquad&P_{2}(U)&=&\frac{1}{2}(3\tanh^{2}U-1)\,,\\[6.0pt] Q_{1}(U)&=&{1-U\tanh U},&\qquad&Q_{2}(U)&=&\frac{1}{2}\big(2U-3\tanh U-3U\operatorname{sech}^{2}U\big)\,.\end{array}$ (V.4)
-The two SturmLiouville{Sturm-Liouville\;} solutions are shown in FIG. 1 in red and in green.
-The $P(U)$ found above provide us with DM geodesics which sare consistent with (II.15). Their caustic behavior at the zeros of $P$ is manifest in FIG.2.
+$Q_{m}=P_{m}S_{m}$ in (III.3), shown in FIG.2 in green, is a second solution of the Sturm-Liouville equation (II.6).
+It is regular for all $U$, vanishes where the Souriau matrix does, and exhibits VM behaviour outside the Wave zone.
+$P_{m}$ and $Q_{m}$ are independent when the Wronskian does not vanish, $W(P,Q)\neq 0$.
+For $m=1$ and $2$ we have, for example,
+[…]
+The $P(U)$ found above provide us with DM geodesics which are consistent with (II.18). The zeros of $P$ are caustic points as it is manifest in FIG.3.
 Figure 3: For a DM wave in $D=1$ transverse dimension Brinkmann trajectories focus to the caustic points determined by $P_{m}(U_{k})=0$, $k=1,\dots,m$.
-For $m=1$ there is one focal point at the origin, and for $m=2$, we have two of them, determined by $\tanh U_{\pm}=\pm{1}/{\sqrt{3}}$.
-Coming to symmetries, the coefficients of the Carroll generators $P$ and $Q$ in (III.1) are shown in FIGs. 3 and 4.
-(a)                       (b)
-Figure 3: For wave number $\bf m=1$}, the (a) translation generator ${\bf P}_{1}$ [which is also a trajectory] is odd.
-(b) the generator of a Carroll boost, ${\bf Q}_{1}={\bf P_{1}S_{1}}$, is even.
-The only zero of ${\bf P}_{1}$ is at $U=0$.
-(a)                       (b)
-Figure 4: Infinitesmal Carroll generators (V) with wave number $\bf m=2$}.
-(a) For a translation, the transverse component ${\bf P}_{2}$ is even, and has two zeros, at $U_{\mp}$.
-(b) For a boost, ${\bf Q}_{2}={\bf P_{2}S_{2}}$ is odd.
-The globally defined Carroll symmetry generators (III.1) are
+For $m=1$ there is one focal point, at the origin, and for $m=2$, we have two of them, determined by $\tanh U_{\pm}=\pm{1}/{\sqrt{3}}$.
+Turning to symmetries, the coefficients of the globally defined Carroll generators $P$ and $Q$ in (III.1), listed in (V.6) and (V) are depicted in FIGs. 4 and 5.
 $\Theta_{m=1}^{Brink}$ $=$ $h\frac{\partial}{\partial V}+c\Big({\tanh U}\,\frac{\partial}{\partial X}-(\operatorname{sech}^{2}U)X\frac{\partial}{\partial V}\Big)$ (V.6)
@@
 $\thinspace{+}b\left(\frac{1}{4}\big(2U-3\tanh U-3U\operatorname{sech}^{2}U\big)\frac{\partial}{\partial X}-\frac{1}{2}\big(1-3\operatorname{sech}^{2}U+3U\operatorname{sech}^{2}U\tanh U\big)X\frac{\partial}{\partial V}\right).\qquad$
-VI Representation of the symmetry generators
-The symmetry-generating vectors along the trajectories, shown in FIGs.3 and 4, are conveniently viewed in a co-moving tangent space – which is a plane carried along the trajectory.
+(a)                        (b)
+Figure 4: For wave number $\bf m=1$}, the (a) translation generator ${\bf P}_{1}$ [which generates also a trajectory] is odd.
+It has one zero, at $U=0$.
+(b) the generator of a Carroll boost, ${\bf Q}_{1}={\bf P_{1}S_{1}}$, is even and has two, symmetrically positioned zeros.
+(a)                        (b)
+Figure 5: Infinitesimal Carroll generators (V) with wave number $\bf m=2$}.
+(a) For a translation, the transverse component ${\bf P}_{2}$ is even, and has two zeros, at $U_{\mp}$.
+(b) For a boost, ${\bf Q}_{2}={\bf P_{2}S_{2}}$ is odd and has 33 zeros.
+VI Graphic representation of the symmetry generators
+The symmetry-generating vectors along the trajectories shown in FIGs.4 and 5 are conveniently viewed in a co-moving tangent space – which is a plane carried along the trajectory.
 In FIGs.
-58 the trajectory is “hidden” in yellow “blobs” and the coordinates $\xi$ and $\eta$ represent the coefficients of ${\partial}_{X}$ and of ${\partial}_{V}$ of the symmetry generators in (III.1), $\Theta^{Brink}={\bm{\xi}}{\partial}_{X}+\eta{\partial}_{V}$, spelt out in (V.5)-(V).
+69 the trajectory is “hidden” in yellow “blobs” and the coordinates $\xi$ and $\eta$ represent the coefficients of the symmetry generators $\Theta^{Brink}={\bm{\xi}}{\partial}_{X}+\eta{\partial}_{V}$ in (III.1), spelt out in (V.6)-(V).
 Figure 6: Infinitesimal translations $({\bm{\xi}},\eta)$ along a DM Brinkmann trajectory ${\bm{X}}(U)$ with $\bf m=1$}, shown in the co-moving tangent space.
@@
 The shortest red arrow $({\bm{\xi}}=1,{\eta}=0)$ is reached for $\bf U=0$} at the caustic point ${\bm{X}}=0,V=0$, where the boost shifts all trajectories by the same amount.
-Figure 7: Translations act along the $m=2$ DM trajectory by (V) The trajectory ${\bm{X}}(U)$ with $\bf m=2$} is “hidden” in the yellow blob.
+Figure 8: Translations act along the $\bf m=2$} DM trajectory, see (V). The trajectory ${\bm{X}}(U)$ with $\bf m=2$} is “hidden” in the yellow blob.
 According to FIG.8, the translation generator (with $c=1/2$) is, for $U=$-$\infty$, $({\bm{\xi}}=1,\,\eta=0)$.
 FIGs.
-3 and 4 confirm that, consistently with (V.5), the two focal points of the trajectory, at $U_{\mp}$, are left invariant.
+4 and 5 confirm that, consistently with (V.6), the two focal points of the trajectory, at $U_{\mp}$, are left invariant.
 For $U=+\infty$ the “8-shaped” curve returns to $({\bm{\xi}}=1,\,\eta=0)$, where it started from.
 Figure 9: Infinitesimal Carroll boosts in Brinkmann coordinates along a chosen DM trajectory marked by a yellow blob, with wave number $\bf m=2$}.
-Boosts along the $m=2$ trajectory are shown in.
-FIG.8.
+Boosts along the $m=2$ trajectory are shown in FIG.9.
 The longest vector on the left (in brown) is, consistently with (V.6) and (V) with $b=1/2$ and $c=2$, the boost for $U=$-$\infty$, and the longest one on the right (also in brown) is for $U=\infty$.
 The two shortest blue arrows (${\bm{\xi}}=\pm 1,{\eta}=0$) show the boost acting as a translations by $\mp b$ at the caustic points $U_{\mp}$.
-Combining FIGs.
-3 with FIG.5 provides us with the outfolding to 3dimensions, shown in FIG.
-9.
-Figure 9: The translations implemented along the trajectories with $m=1$ and $m=2$.
+Combining FIGs.4 with FIG.6 provides us with the outfolding to 3 dimensions, shown in FIG.10.
+Figure 10: Translations implemented along the trajectories with wave numbers $m=1$ and $m=2$.
 The translations implemented along the trajectories vanish where $P(U)=0$. They return to their initial value for $m$ odd and their transverse component changes sign for $m$ even.
-Figure 10: Carroll boosts implemented along the trajectories with $m=1$ and $m=2$.
-Carroll boosts implemented along the trajectories behave at pure translations at P(U)=0P(U)~=~0.
+Figure 11: Carroll boosts implemented along the trajectories with wave number $m=1$ and $m=2$.
+Carroll boosts implemented along the trajectories behave as pure translations at P(U)=0P(U)~=~0.
 They return to their initial value for $m$ odd and their transverse component changes sign for $m$ even.
@@
 VII Conclusion
-Our results shed further light on the intimate relation of Carroll symmetry and the Memory Effect : the matrix $P$ with initial conditions (II.16) yields our geodesics.
-The general expression (III.5) which involves both DM and VM is one of our main results.
+Our results shed further light on the intimate relation of Carroll symmetry and the Memory Effect : the matrix $P$ with initial conditions (II.19) yields our geodesics.
+The general expression (III.6) which involves both DM and VM is one of our main results.
 When the parameter $m$ is an integer, then the wavezone accommodates $m$ half-waves and we get DM.
 For $m=2\ell+1$ $P(U)$ is odd and we do get net displacement.
-For $m=2\ell$, though, $P(U)$ is even and the particle returns, after non-trivial motion in the wave zone, to its initial position : we get DM with no final displacement, consistently with (II.15) and confirmed by FIGs 2, 3a and 4a.
-DM trajectories are distinguished by having zero momentum and can be found by putting ${\bm{p}}_{0}=0$ into (III.5). Turning on the momentum, ${\bm{p}}_{0}\neq 0$, distorts the trajectory to VM.
-The merit of BJR coordinates is their simplicity Sou73 ; Carroll4GW .
-For our initial conditions (II.16), the “motion” reduces to a fixed point, given by the conserved boost momentum ${\bm{x}}_{0}={\bf k}$ in (II.10).
-The entire dynamics is carried by $P(u)$.
-The price to pay is that the BJR description works only between zeros of the $P$-matrix, and must then be fitted together.
-These same points, distinguished by the vanishing of $\det(P)$.
-The advantage of the Brinkmann form (III.1) is that it replaces the necessarily singular Souriau matrix $S$ Sou73 by the globally defined Sturm-Liouville solution $Q$ – another one of our principal results,
-Examplified by the Pöschl - Teller profile, it allowed us to extend the locally given Carrollian Killing vectors to the whole spacetime.
-The clue is that the multiplying by $P$ cures the singularity of the Souriau matrix $S$ and thus that of the BJR coordinate system.
-The Brinkmann trajectories are focused at the zeros of $\det(P)$, as shown in FIG.2.
-The Carroll symmetry generators, first identified by Souriau Sou73 , look more complicated in Brinkmann coordinates as their BJR counterparts do.
-They are however defined globally.
-They are also directly related to geodesics and allow us to unify the Displacement and Velocity Memory effects by making use of the two independent solutions $P$ and $Q$ of the Sturm-Liouville problem, see (III.5).
-The intimate relationship of geodesics with Carroll symmetry is highlighted by their common use of the Sturm-Liouville equation EZHRev ; GenRelGrav .
+For $m=2\ell$, though, $P(U)$ is even and the particle returns, after non-trivial motion in the wave zone, to its initial position : we get DM with no final displacement, consistently with (II.18) and confirmed by FIGs.3, 4a and 5a.
+The Brinkmann trajectories are focused at the zeros of $\det(P)$.
+DM trajectories have zero momentum and zero non-relativistic energy DM-1 .
+They can be found by putting ${\bm{p}}_{0}=0$ into (III.6).
+Turning on the momentum, ${\bm{p}}_{0}\neq 0$, converts the trajectory to VM.
+Both BJR and Brinkmann coordinates have their advantages and drawbacks.
+The merit of BJR is simplicity Sou73 ; Carroll4GW .
+Our initial conditions (II.19) imply ${\bm{p}}=0$, and the “motion” reduces merely to a fixed point determined by the conserved boost momentum ${\bf k}={\bm{x}}_{0}$ and $v_{0}$ — the initial position (II.13)-(II.17).
+The dynamics is carried by $P(u)$.
+The price to pay is that the BJR description is valid only between the mandatory zeros of $\det(P)$, where the $S$-matrix diverges Sou73 ; LongMemory ; EZHRev , and solutions in the adjacent domains have to be fitted together EZHRev .
+The advantage of the Brinkmann form (II.3)–(II.5) is that it replaces the singular Souriau matrix $S$ in (II.10) by the globally defined Sturm-Liouville solutions $P$ and $Q=PS$, (II.6) and (III.3), respectively.
+For zero momentum, (II.16), the non-DM solution $Q$ is turned off, leaving us with $P$ alone which plays its second, trajectory-role.
+The drawback is that finding those solutions may be laborious QLdPT ; DM-1 ; Sila-PLB .
+In conclusion, the main difficulty boils down, in both coordinate systems, to solving a Sturm-Liouville equation.
+Switching to Brinkmann coordinates allowed us to extend the locally given Carrollian Killing vectors Sou73 to the whole spacetime.
+These vectors are directly related to geodesics and allow us to unify the Displacement and Velocity Memory effects by combining the two independent solutions $P$ and $Q$ of the Sturm-Liouville problem, see (III.6), highlighting the intimate relationship of geodesics with Carroll symmetry EZHRev ; GenRelGrav .
+The Brinkmann and BJR forms have also been compared by Harte Harte from a different point of view.
+Showing that the BJR plane wave metric is a generalisation of perturbative weak gravitational waves, he discussed memory effects in a perturbative manner.
 Figs. 611 illustrate the curious behavior of the Carroll generators along the trajectories.
 Further insight is provided by the Schwarzian framework Ovsienko ; ZZH ; Vlahakis .
-This approach confirms that $Q=PS$ is indeed a global, independent, 2nd solution of the Sturm-Liouville problem which does not involve the locally defined Souriau matrix $S$.
-The supersymmetric extension MSUSY sheds further light on the Memory Effect.
-Our results go beyond $D=1$ dimensional profiles and the Pöschl - Teller example.
-Similar results hold for the Scarf profile in $D=2$, which yields an analytic approximation of flyby ZelPol ; Sila-PLB .
-Details will be presented in a forthcoming comprehensive review ZEHPR .
+This approach confirms that $Q=PS$ is indeed a globally defined, 2nd solution of the Sturm-Liouville problem which does not involve the singular Souriau matrix $S$.
+Our results go beyond the toy dimension $D=1$ dimensional profiles and the Pöschl - Teller example.
+Similar results hold for the Scarf profile in $D=2$, which provides us with an analytic approximation of the physically relevant flyby case ZelPol ; Sila-PLB .
+Gravitational waves and their observable effects are also formulated by means of Bondi metric Winicour which relates the gravitational emission and the mass loss.
+The Bondi metric is axially symmetric and asymptotically flat.
+In general, our plane waves are not asymptotically flat, though.
+However, both metrics share similar properties like having a covariantly constant null Killing vector which determines the direction of wave propagation.
+The transverse part of the Bondi metric is quadratic in the radial coordinate which defines the luminosity distance.
+In Harte , the BJR plane wave metric is used to compute similar physical properties.
+To clarify the differences and the similarities between the Bondi formalism with ours here requires more effort and will be reported elsewhere.
+Supersymmetry MSUSY sheds further light on the Memory Effect. Further details will be presented in a comprehensive review ZEHPR .
 Acknowledgements.